ATC $25 Auc $20 HFC-2 FC-? _To if I produce 10 units Arc=25 AVC 2o
Answer #1

Ans: AFC = $5 and FC = $50

Explanation:

ATC = AFC + AVC

$25 = AFC + $20

AFC = $25 -$20 = $5

FC = AFC * Q

= $5 * 10 = $50